OpenStudy (anonymous):

Larry leaves home at 9:03 and runs at constant speed to the lamppost. He reaches the lamppost at 9:07, immediately turns, and runs to the tree. Larry arrives at the tree at 9:13. The lamppost is 200 yards from the start. The tree is 1000 yards from the lamppost. 1. What is Larry’s average velocity, in yards/min, during each of these two intervals. Express your answer using three significant figures. Enter your answers numerically separated by a comma. 2. What is Larry’s average velocity for the entire run? Express your answer using three significant figures.

OpenStudy (anonymous):

Average Velocity...\[\bar{v}=\frac{difference~in~distance~(yards)}{difference~in~time~(minute)}\]From home to lamppost, distance=200 yards, time difference between 9:07 and 9:03 is 7-3 or 4 minutes, therefore from home to lamppost, the average velocity is\[\bar{v}=\frac{200~yards}{4~minutes}=50\frac{yards}{min.}\]

OpenStudy (anonymous):

you can solve the remaining with same approach or analysis... for #2... just get the total distance from home to lamppost to tree... same with the time...